OpenStudy (anonymous):

Change the equation –x + 3y = 9 into slope-intercept form

OpenStudy (anonymous):

The ultimate goal of the problem is to have y on the left-hand side, by itself, and some function of x on the right. What might be the first step in achieving this?

OpenStudy (anonymous):

add 3 to both sides?

OpenStudy (anonymous):

I mean subtract

OpenStudy (anonymous):

Well, the 9 on the right-hand side isn't bothersome, so we can leave it alone. However, if we add x to both sides, the one on the left will disappear, and there will be +x on the right.

OpenStudy (anonymous):

Slope intercept form is of the form y=mx+c; where m is the slope and c is the y-axis intercept. Separate the y term to one side of the equation and then divide both sides by the coefficient(if any) of the y term. Example: 2x+9y-13=0 in the slope intercept form will be 9y=-2x+13 and finally y=(-2/9)x+(13/9) which is of the form y=(m)x+(c). Now solve your equation accordingly.

OpenStudy (anonymous):